# Can Robots “Taste” Grapes? Estimating SSC with Simple RGB Sensors

### v1

@@ 2.1.2 Data collection and labeling
We collected static images during two summer seasons using different camera sensors. Three different devices, across multiple years, have been used to collect the images. We employed a smartphone, a Reflex, and a Realsense d435i to represent some of the possible covariate shifts that occur between the collection of the initial data on the field by a farmer or an agronomist (smartphone), a high-quality sensor (reflex) and a low-quality robotic camera (d435i). Table 1 depicts the working categories we used in this study. A more detailed description of all the data collected and the different characteristics is given in Appendix A for the interested reader.
@@
Name Split # samples SSC mean SSC std SSC MAD Description
$T1_{STD}$ 112 17.19 2.43 1.99 Phone images in JPG format captured in the 2021 season
$T1_{WB}$ 112 17.19 2.43 1.99 Phone images color corrected from RAW captured in the 2021 season
$T2_{STD}$ tr 212 16.23 2.52 2.02 Phone images captured in JPG format during the 2021 season.
$T2_{STD}$ te 92 18.91 2.16 1.71 Phone images captured in JPG format during the 2022 season.
$T2_{WB}$ tr 150 17.54 2.46 2.04 Phone images captured in RAW format during the 2021 season.
$T2_{WB}$ te 92 18.91 2.16 1.71 Phone images captured in RAW format during the 2022 season.
$T2_{ALL}$ tr 300 17.54 2.46 2.04 Phone and reflex images in RAW format, 2021 season.
$T2_{ALL}$ te 92 18.91 2.16 1.71 Phone and reflex images in RAW format, 2022 season.
$T3_{STD}$ tr 352 17.67 2.75 2.33 All jpg images, across all seasons and devices.
$T3_{STD}$ te 44 15.96 1.68 1.36 The test set contains images with SSC content close to the threshold (16° Brix).
$T3_{WB}$ tr 440 18.27 1.98 1.96 All white balanced images, across all seasons and devices.
$T3_{WB}$ te 44 15.96 1.36 1.36 The test set contains images with SSC content close to the threshold (16° Brix).
Table 1: $T1_{STD}$ and $T1_{WB}$ consist respectively in jpg and raw versions of images collected with a smartphone camera during the 2021 season. $T2$ datasets are images collected with a smartphone camera, but where the training splits are from the 2021 season, and the test set from the 2022 season. In this case, there is no exact correspondence between $T2_{STD}$ and $T2_{WB}$ jpg and raw images. The $T3$ datasets serve a different purpose, as they combine images from the $T1$ and $T2$ datasets, mixing data across all seasons and devices, specifically the smartphone, Reflex (only for the WB variant) and d435i cameras. The $T3$ training set has more samples than the $T2$. At the same time, the $T3$ test set consists of images with SSC threshold quite close to the ripeness level for harvesting (16° for Pizzutello), making it significantly more challenging than the training set.
@@
Labels: All the images were labeled both for SSC and color. For SSC, we used an Atago Hikari PAL2 digital refractometer, capable of measuring berry sugar content (degree Brix, °Brix) by contact without damaging the berry itself. This instrument has a range from 9° to 25°. Typically, depending on the grape variety, an SSC of 16°-18° is considered good enough for harvesting. Since we are trying to assess how well the estimation based on appearance would perform compared to human performance, we adopted a labeling procedure similar to the SSC sampling performed by agronomists. Figure 2 visually illustrates the process. Every berry in a grape bunch has a different SSC, and the sugar concentration in each berry is also not uniform due to gravity. For this reason, common practice is to sample the SSC three times, from the top, middle, and bottom of the bunch, and then average the results. For consistency, the refractometer was placed on the side of the berry in each measurement. For color, a categorical classification was required. Based on agronomic production use cases, we defined a color quality chart with the help of experienced agronomists. The color scale goes from 1 to 6, with whole number increments, where 1 represents the best quality and 6 the worst. Usually, values of 1 and 2 are considered high-quality produce, while 3 is still acceptable for some markets. All the other values are considered unripe. This color index accounts for color saturation and distribution across the berries and the bunch. An example of the color variability is given in Fig: 3.
@@
Figure 3: Examples of the six color levels for color classification. A clear distinction between categories may be challenging for an untrained eye. However, our labeling validation tests described in Section 3.1.2 indicate consistent labeling despite some uncertainty in the classification of neighboring classes.
2.2 Estimation with Histogram Features
@@
The main use-case for the estimators we design in this work is to be deployed on a robotic platform, embedded edge devices, or smartphone-like device, thus with limited memory and computational power. For this reason, the first approach is based on classic hand-crafted features, specifically on histograms. RGB channel histograms were chosen as the basic building blocks since the problem largely depends on color intensity and distribution. To capture distribution differences, we divided each bunch image into a grid of rectangular regions and computed the RGB histogram for each region. A key observation that can be made with this setup is that even with tight bounding boxes from the detector, some regions will often contain background information due to the shape of grape bunches. To address this issue, we modified the feature processing to retain only specific regions in a ”cross-shaped” selection, as illustrated in Figure 4. This selection focuses on the areas of the image where grape content is most likely to be present.
Figure 4: Each grape bunch in the scene is cropped based on the bounding box provided by the detector. The resulting image is then binned, with some bins dropped as they contain background information with high probability. This picture shows an example, but the actual bin size and dropping schemes are hyper-parameters described in Section 2. In this simple example, the n-uple defining column and row ranges described by Eq.1 would be [1,3,1,2]
To simplify hyper-parameter exploration, we defined three approaches for region selection: a ”fat” cross-shaped pattern, a ”thin” one, and one using all regions without any exclusion. More in detail, the three configurations are defined by the starting and ending cell in the column and row direction as follows:
@@
We compared these approaches to determine the most effective feature selection strategy. In summary, the hyper-parameters that were considered while designing the features are:
  • 1.
    Image starting size (width, height)
  • 2.
    Number of cells in both the width and height directions
  • 3.
    The cross pattern used to remove the background
  • 4.
    Use of HSV instead of RGB channels
We performed a hyper-parameter search for both the task of SSC and color degree estimation using Ridge Regression. For brevity, since color estimation is an easier task to perform, in Section 3 we omit the details of the experiments for color estimation, while the SSC ones are explored in depth.
2.3 Estimation with Multi-Task Learning DNN Architecture
As mentioned before, one of the research questions we are exploring is the correlation between the color of the berries and sugar content. The correlating latent factors are many and difficult to model explicitly. Among them are climate, weather, sunlight exposure, pruning strategies, and the use of nitrogen and other chemicals. Given this complex interaction between different factors, we decided to design a neural architecture that could extract common features to be used for both estimation tasks.
@@ 2.3 Estimation with Multi-Task Learning DNN Architecture
Our model is shown in Figure 5. To distill a feature vector, we employed an Auto-Encoder (AE) architecture, instead of using only a feature extraction backbone alone. The reason is that we are in a data-scarcity scenario, so having the additional reconstruction loss of the auto-encoder, focusing on appearance, helps in strengthening the back-propagation signal and reduces over-fitting. Our auto-encoder is a UNet with a ResNet18 backbone. In general, UNet-like architectures alone are not suitable for reconstruction problems, since the skip connections would cause the network to simply learn an identity function. Our architecture avoids this problem, by using the bottleneck feature output as the input for the feature extraction section. From those features, a convolutional block is used to extract the features that are then fed to the task-specific heads. The two heads have the same structure, but in the case of color estimation, the outputs are logits for the multi-class problem, while for the SSC estimation are the actual estimations.
@@
The loss for the architecture is the combination of the reconstruction loss of the AE (pixel-wise Mean Squared Error (MSE)), with the addition of a Mean Absolute Error (MAE) loss for the SSC estimate, and a Cross-Entropy loss for the color classes:
@@
Figure 5: The multi-task network is based on a U-Net [13] with a ResNet18 [7] backbone, designed to encode the image into a compact representation. Features extracted from the bottleneck layer are fed into a two-head estimator: one for SSC (regression) and another for color (classification). The two heads share a common ”neck,” a CNN with 512 output channels, ensuring part of the representation is shared between the tasks. Each head has a symmetrical structure, comprising three CNN blocks that progressively reduce the feature size, followed by adaptive average pooling [24] and a 1x1 convolution [10] for the final estimation.
2.3.1 Hyper-parameter Validation
Our evaluation criteria also align with agronomic quality standards, emphasizing the importance of avoiding the harvest of unripe bunches, while allowing for some variation in quality among ripe bunches. This is achieved by introducing hinge-like penalties that apply only when estimations fall below predefined quality thresholds for SSC ($t_{brix}$) and color ($t_{color}$) estimation, ensuring that models are primarily evaluated on their ability to meet minimum quality standards. The final score, $\mathcal{S}$ integrates the normalized performance metrics and the agronomic quality criteria.
The formal definition of the score $\mathcal{S}$ is the following:
$L_{BH}$ $=\max(0,L_{brix}-t_{brix}))^{\rho}$ (4)
$L_{CH}$ $=\max(0,-\eta(A_{color}-t_{color}))^{\rho}$ (5)
where $\rho$ is the degree of the two hinge penalties $\mathcal{L}_{BH}$ (for °Brix) and $\mathcal{L}_{CH}$ (for color), set to 2 in all our experiments. $t_{brix}$ and $t_{color}$ represent the agronomic quality thresholds for °Brix and color, respectively, and $\eta$ is the normalization factor between SSC and color scores. In our experiments, $t_{brix}=2.5$Brix, which is comparable with the standard deviation of the training data, and $t_{color}=0.8$, since an error rate of $80\%$ is usually considered acceptable for human performance in this context [14]. The normalization factor is empirically set as
$\eta=\frac{t_{brix}}{t_{color}}.$ (6)
@@ 3 Experiments
In designing the experimental setup, we had two main aims. The first one was to give an initial answer on factors are relevant in estimating SSC from appearance, in order to better understand the problem for future work and data collection campaigns. The second aim is to validate the proposed features and estimator models. This section begins with the validation of the collected dataset, followed by a description of the exploratory experiments. Finally, we describe the validation of our proposed features and models, together with the ablation studies. The data and code for all the experiments can be found on the accompanying GitHub repository [^1] To be disclosed upon publication.
3.1 Exploratory User Studies
@@ 3.1.1 Human Evaluated SSC and Color Correlation
Feature description
v1, v2, v3 Upper, middle and lower °Brix measurements.
m_brix Mean °Brix computed from v1, v2, v3.
m_b_t Rounded m_brix (unused).
b_class Binary °Brix class for harvesting (Sufficient, Insufficient). Computed from m_brix with threshold $17$°.
c_val Expert given score for color development after varaison, on a scale from 1 to 10.
c_class Binary color class for harvesting (Sufficient, Insufficient). Computed from c_val with threshold $6$.
l_val, r_val Expert given scores for lignification and berry shape development, respectively, on a scale from 1 to 10.
Table 2: These are the parameters considered by the agronomist in the exploratory experiment to assess the correlation between °Brix and human color evaluation.
@@
In order to have an initial human performance baseline and to assess the experts’ actual capacity to estimate ripeness from appearance, we performed a specific test. We gave to a number of agronomists with experience in table grape ripeness estimation, $100$ grape images, for which a °Brix ground truth was collected with the technique described in Fig. 2. We used the JPG smartphone images collected during the first data collection session (24 September 2021) for Pizzutello Nero. The images were given without any other hint to the experts, asking them to express a numeric evaluation, on a scale from 1 to 10, of the degree of color, peduncle lignification stage, and berry shape. Note that this color scale is not the same as the one defined in Section 2.1.2. The reason is that this experiment uses a human-friendly decimal scale, while the one defined previously is to be used by a classification algorithm. In addition to the aforementioned parameters, the agronomists were asked to guess if the °Brix for the sample was high enough for harvesting purposes (in this case, $\geq 17$). Table 2 summarizes the experiment variables name and meaning. First, we checked the correlation between human-given values of c_val, l_val and r_val and the measured mean °Brix value, m_brix to explore the hypothesis where color and other visual elements could be a good predictor for sugar content.
@@
In Fig. 6 the correlation matrix is shown. The correlation score between the average °Brix (m_brix) and color evaluation (c_val) is $0.505$, with $p<0.001$ and $CI=[0.342,0.638]$, the correlation between m_brix and lignification (l_val) is $0.337$, $p<0.001$ and $CI=[0.151,0.501]$, while the correlation with berry shape (r_val) score is much lower, $-0.038$ way below any significance threshold.
@@
The observed correlation between human expert visual assessment and measured °Brix values suggests that a data-driven algorithm could achieve human-level accuracy in estimating grape quality for harvesting decisions. Even if there is a significant correlation with the lignification of the peduncle, in this work we focused on the color, leaving the use of peduncle lignification as a predictor for future work.
The second outcome of this experiment was to define some human harvesting performance baselines. For this question, we consider the c_class, the color category given by the agronomists to each sample, as the harvesting prediction and the thresholded values b_class as the binary labels. The precision and recall values are respectively $0.43$ and $0.83$, which means a false discovery rate (type I error rate) of $0.56$ and false negative rate (type II error rate) of $0.17$.
@@
It is important to state that in the case of harvesting, type I errors (false positives) have a higher cost than type II since all the ripe bunches left on the vine tree will be eventually harvested, while the harvested unripe bunches have to be disposed of. We see that by simply collecting the table grape bunches with a color class over the threshold of $6$ (the definition for the c_class attribute) a lot of unripe bunches are collected due to °Brix to color mismatch. One question that arises then is if by increasing the threshold for c_class, it would be possible to trade type I errors for type II errors. Computing the precision and recall scores for a c_class value obtained with a threshold of $7$ for example, we get respectively $0.35$ precision and $0.37$ recall, which is an increase in both error types. This test shows that, while °Brix and color are correlated on the broad range of the color scale of this experiment, at the same time, the relationship is not quite linear, i.e. there are both a considerable number of insufficiently colored bunches that have sufficient °Brix, and highly colored bunches with insufficient °Brix. The high number of bunches on a borderline situation that the experts miss-classify shows that the question of training an appearance-based estimator is a non-trivial one.
Figure 6: Correlation matrix of the parameters estimated by human experts looking to a sample of 50 images of table grape bunches, as described in Section 3.1. The key factor for our evaluation is the correlation between the SSC guess (m_brix) and the evaluation of bunch color (c_val).

### v2

@@ 2.1.2 Data collection and labeling
We collected static images during two summer seasons using different camera sensors. Three different devices, across multiple years, have been used to collect the images to represent some of the possible covariate shifts that occur between the collection of the initial data on the field by a farmer or an agronomist and the deployment on a robotic hardware.
Images capture: We acquired images using two different smartphones, a Reflex camera, and a Realsense d435i. The smartphone images were captured with a Moto G8 Plus and a Realme GT Master Edition, respectively at $3000\times 4000$ and $3448\times 4624$ pixel resolution in JPG and raw formats. The Reflex images were captured in raw format at a $3900\times 2613$ pixel resolution, while the d435i provided RGB images at a lower $1280\times 720$ pixel resolution. When available, raw images were corrected for white balance using a gray card (Fig. 3), though images without a gray card were still used for comparison. Fig. 4 shows some examples of images collected with the same smartphone both in JPG and white-balanced RAW format. All images were labeled with bounding boxes around the grape bunches and subsequently cropped. The resulting images vary in size, as some are close-ups while others contain multiple targets. The combination of season, sensor, and gray card presence creates differences summarized in Table 4. The rows in the table represent atomic splits—groups with the same data distribution—and the variations between them introduce the covariate shifts we aim to study.
Dataset Id Year Smartphone Reflex d435i Num images Num samples Expert GT Color GT Type Jpg Raw Card Raw Card RGB IR Card A 2021 motog8 yes no no no - no no - 100 100 no no B 2021 motog8 no yes yes yes yes no no - 38 38 possible yes C 2021 motog8 yes yes yes yes yes no no - 62 62 possible yes D 2021 motog8 yes yes yes yes yes no no - 50 50 possible yes E 2022 realmi yes yes no no - yes no no 1 2 no no F 2022 realmi yes yes yes no - yes no yes 6 22 possible yes G 2022 realmi yes yes no no - yes no yes 1 2 no no H 2022 realmi yes yes yes no - yes no yes 4 19 possible yes I 2022 motog8 yes yes yes no - yes yes yes 8 51 possible yes
@@
Figure 4: Some examples of paired JPG and RAW (white-balanced) images collected with the same phone device.
@@
Labeling protocol: All the images were labeled both for SSC and color. For SSC, we used an Atago Hikari PAL2 digital refractometer, capable of measuring berry sugar content in °Brix by contact without damaging the berry itself. This instrument has a range from 9° to 26°. Typically, depending on the grape variety, an SSC of 16°-18° is considered high enough for harvesting. Since we are trying to assess how well the estimation based on appearance would perform compared to human performance, we adopted a labeling procedure similar to the SSC sampling performed by agronomists. Fig. 3 visually illustrates the process. Every berry in a grape bunch has a different SSC, and the sugar concentration in each berry is also not uniform due to gravity. For this reason, common practice is to sample the SSC three times, from the top, middle, and bottom of the bunch, and then average the results. For consistency, the refractometer was placed on the side of the berry in each measurement. For color, a categorical classification was required. Based on agronomic production use cases, we defined a color quality chart with the help of experienced agronomists. The color scale goes from 1 to 6, with whole number increments, where 1 represents the best quality and 6 the worst. Usually, values of 1 and 2 are considered high-quality produce, while 3 is still acceptable for some markets. All the other values are considered unripe. This color index accounts for color saturation and distribution across the berries and the bunch. An example of the color variability is given in Fig. 5.
@@
Figure 5: Examples of the six color levels for color classification. A clear distinction between categories may be challenging for an untrained eye. Agronomist classify the color category taking into account both the general color shades of the bunch and the density of highly colored berries. However, our labeling validation tests described in Section 3.1.2 indicate consistent labeling despite some uncertainty in the classification of neighboring classes.
Experimental Dataset Splits: Our dataset is built upon a foundation of atomic splits, which are minimal, homogeneous groups of images categorized by season, device, and processing type, as detailed in Table 4. We combined these atomic splits to create several experimental datasets designed to evaluate model performance in specific scenarios. These combined datasets are distinguished by their image type: standard JPG images (indicated by the STD subscript) or white-balanced images derived from raw captures (indicated by the WB subscript). A summary of these experimental datasets is presented in Table 5.
The datasets are grouped by their analytical purpose:
  • 1.
    The Simple splits are small, single-season (2021) datasets used for initial, exploratory assessments of our models. They do not have a predefined train/test division.
  • 2.
    The Season21ToSeason22Phone splits fix the sensor type to smartphones and are designed to evaluate cross-seasonal performance by training on data from one season and testing on the next.
  • 3.
    The MixedSeason splits represent the most challenging scenario. They are designed to simulate a real-world deployment by training on smartphone and Reflex images (mimicking data collection by agronomists) and testing on d435i images (mimicking a robotic platform). This directly tests the models’ robustness to the covariate shift induced by different acquisition devices.
To provide a quantitative baseline for our experiments, Table 6 summarizes key statistical properties of these datasets. The table details the size of each train and test split, along with the natural variability of the SSC labels, measured by the mean, standard deviation, and Mean Absolute Deviation (MAD). These statistics will serve as a reference for evaluating the performance of our proposed estimation models.
Table 5: Summary of the experimental dataset splits used for model training and evaluation.
[…]
Table 6: Statistical properties of the experimental dataset splits.
[…]
2.2 SSC Estimation Models
In this Section, the two algorithmic approaches to SSC estimation are described.
2.2.1 Estimation with Histogram Features
@@
The main use-case for the estimators we design in this work is to be deployed on a robotic platform, embedded edge devices, or smartphone-like devices, thus with limited memory and computational power. For this reason, the first approach is based on classic hand-crafted features, specifically on histograms. RGB channel histograms were chosen as the basic building blocks since the problem largely depends on color intensity and distribution. To capture distribution differences, we divided each bunch image into a grid of rectangular regions and computed the RGB histogram for each region. A key observation that can be made with this setup is that even with tight bounding boxes from the detector, some regions will often contain background information due to the shape of grape bunches. Adding a background subtraction deep network, such as in Qin et al., [14], could have been effective to some extent, at the cost of changing the scale of the computational costs. For this reason, we decided to opt for a more straightforward approach. To address this, we developed a simple heuristic to remove regions likely to contain background pixels. Based on the empirical observation that grape bunches are typically wider in the middle and narrower at the top and bottom, we retain only a ”cross-shaped” selection of regions, as illustrated in Fig. 6. We treat the exact dimensions of this cross-shaped pattern as a hyper-parameter to be optimized. To simplify this exploration, we defined three configurations in Equation 1: none (keeping all regions), fat (dropping corner regions), and thin (dropping corners and more bottom regions).
More in detail, the three configurations are automatically defined by the selection of the number of horizontal ($n_{bin_{x}}$) and vertical ($n_{bin_{y}}$) divisions (the width and height of the grid in Fig. 6) as follows:
@@
Averaging the RGB values on each cell and concatenating them produces a final feature vector with length of $3\times n_{bin_{x}}\times n_{bin_{y}}$. In our experimental validation, We considered also the exploration of the effects of the initial image size and the use of Hue-Saturation-Value (HSV) color space. Given the small number of images, to avoid overfitting, we also considered $L2$ regularization. We use Ridge Regression as the estimation algorithm, since it is the simplest interpretable regression algorithm that allows for regularization. Clearly, other more sophisticated models – such as Support Vector Machines of Random Forests – can be used for better estimation results, but in this work we are interested in exploring the characteristics of the features and not only the peak performances.
Figure 6: An illustration of the ”cross-shaped” feature selection process for histogram-based estimation. An input image is divided into a grid of $n_{bin_{x}}$ by $n_{bin_{y}}$ bins ($5\times 5$ grid shown). To heuristically remove background pixels, two patterns are tested: a fat pattern that retains a wider central cross-section and a thin pattern that is narrower at the bottom, mimicking the typical shape of a grape bunch. The indices of the kept bins are determined empirically using Equation 1. For the $5\times 5$ example shown, this results in the index ranges [1,3,1,3] for the fat pattern and [1,3,1,2] for the thin pattern.
2.2.2 Estimation with Multi-Task Learning DNN Architecture
The SSC and color distribution can be viewed as conditioned on the environmental and agronomic parameters of the field we are considering. While both the color and the SSC depend on the internal chemical state of the berries, this state itself is influenced by macroscopic environmental and managerial factors such as growing conditions, training practices, and nutrient availability. The overall appearance of the grape bunches is also affected by these factors, which we assume largely homogeneous with respect to key management practices and environmental conditions. Nevertheless, our proposed methods are designed to be robust to the inherent noise and micro-variations arising from individual plant differences, which are expected in any real-world field data.
@@ 2.2.2 Estimation with Multi-Task Learning DNN Architecture
The architecture of the multi-task network, illustrated in Fig. 7, reflects this reasoning. To distill a feature vector, we employed an Auto-Encoder (AE) architecture, instead of using only a feature extraction backbone alone. The encoder component captures common features in appearance variations that may correlate with the internal chemical state of the grapes. These shared features are then leveraged by two separate prediction heads to estimate the color class and SSC in parallel. The reason is that we are in a data-scarcity scenario, so having the additional reconstruction loss of the auto-encoder, focusing on appearance, helps in strengthening the back-propagation signal and reduces over-fitting. Our auto-encoder is a UNet [15] with a ResNet18 [7] backbone. The two heads share a common ”neck,” a CNN with 512 output channels, ensuring part of the representation is shared between the tasks. Each head has a symmetrical structure, comprising three CNN blocks that progressively reduce the feature size, followed by adaptive average pooling Zhou et al., [27] and a 1x1 convolution Lin et al., [11] for the final estimation. In general, UNet-like architectures alone are not suitable for reconstruction problems, since the skip connections would cause the network to simply learn an identity function. Our architecture avoids this problem, by using the bottleneck feature output as the input for the feature extraction section. From those features, a convolutional block is used to extract the features that are then fed to the task-specific heads. The two heads have the same structure, but in the case of color estimation, the outputs are logits for the multi-class problem, while for the SSC estimation are the actual estimations.
Figure 7: The multi-task network architecture used for SSC and color estimation. The model is based on a U-Net (Ronneberger et al., [15]) with a ResNet18 (He et al., [7]) backbone, which acts as an auto-encoder to learn a compact image representation (outputting the reconstruction loss, $\mathcal{L}_{rec}$). Features from the Bottleneck Conv layer are passed to a block of Shared features. From here, the network splits into two parallel heads: a Brix regressor for SSC estimation (outputting $\mathcal{L}_{brix}$) and a Color classifier for quality grading (outputting $\mathcal{L}_{color}$). (Diagram generated with PlotNeuralNet [8]).
@@
The loss for the architecture is a classical compound loss that combines the reconstruction loss of the AE (pixel-wise MSE), with the addition of a Mean Absolute Error (MAE) loss for the SSC estimate, and a Cross-Entropy loss for the color classes:
@@
To optimize the hyper-parameters of our multi-task DNN, we developed a custom scoring metric designed to balance statistical accuracy with agronomic needs. We call this the Brix-Color-Hinge (BCH) metric, as it integrates the regression error for °Brix, the classification accuracy for color, and includes hinge-like penalties for predictions that fall below minimum quality thresholds. This approach ensures that the selected model is not only accurate but also practical for harvesting decisions. The detailed mathematical formulation of the BCH metric is provided in C.
@@ 3 Experiments
This Section describes the experiments of the two proposed algorithmic solutions to SSC estimation. In order to give perspective to these results, some experiments were performed to set a human performance baseline. For this reason, the Section starts with the explanation on how human performances were tested in this scenario (Section 3.1), then proceeds with the actual experiments of the histogram feature model (Section 3.2) and multi-task DNN model (3.3). Both algorithmic solutions have been initially developed and trained locally on a laptop equipped with a AMD Ryzen 9 5980HX Mobile CPU and a NVIDIA GeForce RTX 3080 GPU with 16GB of VRAM, running Ubuntu 22.04, while the hyper-parameter optimization for all models have been run on a NVIDIA DGX workstation equipped with A100 GPUs. All the code has been written in python3. The data and code for all the experiments can be found on the accompanying GitHub repository [^1] https://github.com/Smart-Agri-DIAG/robots-tasting-grapes.
3.1 Human Performance Baselines
@@ 3.1.1 Human Evaluated SSC and Color Correlation
To establish an initial human performance baseline and assess experts’ ability to estimate ripeness from appearance, we performed a specific test. We presented $100$ grape images to a number of agronomists with experience in table grape ripeness estimation. For these images, a °Brix ground truth was collected with the technique described in Fig. 3. We used the JPG smartphone images collected during the first data collection session (24 September 2021) for Pizzutello Nero. The images were presented without further context, and the experts were asked to assign a numerical value, on a scale from 1-to-10, to the degree of color, peduncle lignification stage, and berry shape, respectively indicated in the following as color_score, lignification, and berry_shape. Note that this 1-to-10 color scale differs from the 1-to-6 scale defined in Section 2.1.2; This experiment utilized a decimal scale for ease of human interpretation. To create a binary harvesting prediction from this data, we used the agronomists’ color_score as a proxy for their final decision, applying a threshold of 6. We called the thresholded prediction harvesting_pred. We then measured the correlation between all three visual estimates and the measured mean_brix to validate our core hypothesis. The correlation scores are shown in Fig. 8.
@@
The correlation score between the average °Brix (mean_brix) and color evaluation (color_score) is $0.505$, with $p<0.001$ and $CI=[0.342,0.638]$, the correlation between mean_brix and lignification (lignification) is $0.337$, $p<0.001$ and $CI=[0.151,0.501]$, while the correlation with berry shape (berry_shape) score is considerably lower, at $-0.044$, which is below any significance threshold.
@@
The observed correlation between human expert visual color assessment and actual measured °Brix values suggests that a data-driven algorithm could achieve human-level accuracy in estimating grape quality for harvesting decisions.
Figure 8: Correlation matrix of key parameters from the human evaluation study on 100 grape images. The matrix shows the relationship between the measured mean_brix and the scores assigned by experts for color_score, lignification, and berry_shape. The strong correlation of 0.505 between mean_brix and color_score is the key finding that validates using visual appearance for ripeness estimation.
This experiments also helps in setting a human performance baseline. Given that a °Brix value greater than $17$ is considered sufficient for harvesting, we use the thresholded mean_brix as the harvesting ground truth label, comparing it to harvesting_pred. The precision and recall values were $0.43$ and $0.83$, respectively, translating to a false discovery rate (Type I error rate) of $0.57$ and a false negative rate (Type II error rate) of $0.17$. These values will be compared with similar scores computed with the proposed algorithms in the following Sections.
@@
As a final note to this analysis, it is important to note that in the case of harvesting, Type I errors (false positives) carry a higher cost than Type II errors. This is because ripe bunches left on the vine will eventually be harvested, whereas unripe harvested bunches must be discarded. By collecting table grape bunches with a color class above the threshold of $6$ (the definition for the harvesting_pred attribute) a significant number of unripe bunches are collected due to a mismatch between °Brix and color. This raises the question of whether increasing the threshold for harvesting_pred could trade Type I errors for Type II errors. However, computing the precision and recall scores for a harvesting_pred value obtained with a threshold of $7$ yields $0.35$ and $0.37$, respectively, indicating an increase in both error types. This test shows that, while °Brix and color are correlated across the broad range of the color scale used in this experiment, at the same time, the relationship is not strictly linear. There are a considerable number of both insufficiently colored bunches with sufficient °Brix and highly colored bunches with insufficient °Brix. The high number of bunches in a borderline situation that the experts miss-classify shows that training an appearance-based estimator is a non-trivial task.